# Fuzzy K-Means Clustering without Cluster Centroids

### v1

@@ 3.2. Distance matrix
3.2.1. Squared Euclidean distance
The squared Euclidean distance is the most commonly used distance metric. It measures the absolute distance between points in a Euclidean space. The distance between the $i$-th and $j$-th samples can be expressed as follows:
(13) $d_{euc(ij)}=\left\|\textbf{x}_{i}-\textbf{x}_{j}\right\|_{2}^{2}$
3.2.2. K-Nearest Neighbor Distance
The K-nearest neighbor distance refers to the process of computing the distances between a given query point and all other points in the dataset, and then selecting the K nearest points. For instance, in Euclidean clustering, the K-nearest neighbor distance is computed as follows:
(14) $d_{knn(ij)}=\begin{cases}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2},&\textbf{x}_{i}\leftrightarrow\textbf{x}_{j}\\ \sigma,&\text{otherwise}\end{cases}$
@@ 3.2.2. K-Nearest Neighbor Distance
3.2.3. Adjacency Matrix
Given an adjacency matrix $\mathbf{S}$, the distance matrix $\mathbf{D}$ can be computed utilizing a Butterworth filter as proposed in (Lu et al., 2023). The process is outlined as follows:
(15) $d_{adj(ij)}=\sqrt{\frac{1}{1+(\frac{s_{ij}}{\Omega})^{4}}}$
@@ 3.3. Optimization Algorithm
(17) $J(\textbf{Y})=\textrm{tr}(\textbf{Y}^{\textrm{T}}\textbf{D}\textbf{Y}\textbf{P}^{-1})+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$
@@
(18) $\frac{\partial J}{\partial\textbf{Y}}=(\textbf{D}+\textbf{D}^{\textrm{T}})\textbf{YP}^{-1}+2\lambda\textbf{Y}+\frac{\partial(\textbf{a}^{\textrm{T}}\hat{\textbf{p}})}{\partial\textbf{Y}}$
@@
(19) $\frac{\partial(\textbf{a}^{\textrm{T}}\hat{\textbf{p}})}{\partial y_{ij}}=\frac{\partial}{\partial y_{ij}}\left(\sum\limits_{j=1}^{K}{a_{j}p_{jj}^{-1}}\right)=-a_{j}p_{jj}^{-2}$
@@
(20) $\frac{\partial J}{\partial y_{ij}}=g_{ij}-a_{j}p_{jj}^{-2}$
@@
(22) $y_{ij}\leftarrow y_{ij}\sqrt{\frac{a_{j}p_{jj}^{-2}}{g_{ij}}}$
@@
7:     Normalize Y by $y_{ij}=\frac{y_{ij}}{\sum\nolimits_{j}{y_{ij}}}$
@@
3.4. Computational Complexity Analysis
@@ 4. Experiments
4.1. Comparison Algorithms
  1. (1)
    K-Means++ (Arthur and Vassilvitskii, 2007):The advantages of careful seeding for (K)-Means.
  2. (2)
    Ksum (Pei et al., 2023b): This algorithm integrates spectral clustering and K-Means into a unified framework.
  3. (3)
    Ksum-x (Pei et al., 2023b): Distinct from Ksum, Ksum-x inputs a data matrix instead of a K-NN graph.
  4. (4)
    RKM (Lin et al., 2019): This algorithm minimizes the squared distance from the samples to the clustering centers and incorporates a regularization term for balanced clustering.
  5. (5)
    CDKM (Nie et al., 2022a): Demonstrates that the underlying model of K-Means can be transformed into a matrix trace maximization problem, proposing an efficient coordinate descent solution.
  6. (6)
    MSFCM (Zhou et al., 2021)A new membership scaling Fuzzy C-Means clustering algorithm.
  7. (7)
    ULGE (Nie et al., 2017) Unsupervised Large Graph Embedding.
4.2. Benchmark Datasets
The performance of our method is evaluated on seven benchmark datasets, and the information of size is shown in Table 1
  1. (1)
    AR (Martinez and Benavente, 1998) includes 120 kinds of face with 3120 images.
  2. (2)
    JAFFE (Lyons et al., 1998)
  3. (3)
    MSRC_V2 (Winn and Jojic, 2005) which includes 7 kinds of objects with 210 images. We selected the 576-D HOG feature as the single view dataset.
  4. (4)
    ORL (Cai et al., 2010) includes 400 pictures of face from 40 people.
  5. (5)
    UMIST (Hou et al., 2014) consists of 564 images of face from 20 individuals
  6. (6)
    USPS (Hull, 1994) consists of 3000 handwritten digital images for each of the 10 digits from 0 to 9.
  7. (7)
    Yaleface (Belhumeur et al., 1997) includes 165 grayscale images in GIF format of 15 people
Table 1. The information of the benchmark datasets, including
@@ 4.2. Benchmark Datasets
Table 2. ACC results on 7 benchmark datasets from 7 comparative algorithms and FKMWC with distance by Eq. (15) (Ours_bw) ,squared Euclid distance (Ours_d) and K nearest neighbors distance (Ours_kd), with the highest value represented in bold.
@@
Datasets K-Means++ Ksum Ksumx RKM CDKM MSFCM ULGE Ours_bw Ours_d Ours_kd
AR 0.2514 0.2970 0.2454 0.2641 0.2653 0.2747 0.3442 0.3917 0.2843 0.3728
JAFFE 0.7085 0.8789 0.8930 0.8310 0.7108 0.8357 0.9249 0.9155 0.8404 0.9671
MSRC_V2 0.6052 0.6852 0.6852 0.6286 0.6657 0.7381 0.7429 0.7714 0.7667 0.8095
ORL 0.5198 0.6337 0.5877 0.5000 0.5507 0.5725 0.6200 0.6625 0.6075 0.6875
UMIST 0.4339 0.4209 0.4296 0.4209 0.4210 0.4591 0.5043 0.4974 0.4713 0.5252
USPS 0.6491 0.7664 0.7240 0.6673 0.6416 0.6577 0.7530 0.7923 0.7013 0.7553
Yaleface 0.3812 0.4339 0.4418 0.4485 0.3964 0.4545 0.5212 0.4909 0.4788 0.5273
@@
Table 3. NMI results on 7 benchmark datasets from 7 comparative algorithms and FKMWC with distance by Eq. (15) (Ours_bw) ,squared Euclid distance (Ours_d) and K nearest neighbors distance (Ours_kd), with the highest value represented in bold.
@@
Datasets K-Means++ Ksum Ksumx RKM CDKM MSFCM ULGE Ours_bw Ours_d Ours_kd
AR 0.5574 0.5963 0.5676 0.5752 0.5700 0.5811 0.6090 0.6500 0.5722 0.6268
JAFFE 0.8010 0.8764 0.9013 0.8159 0.7981 0.8806 0.9005 0.9225 0.8545 0.9623
MSRC_V2 0.5280 0.5753 0.5753 0.5612 0.5693 0.6010 0.5870 0.6265 0.6475 0.6712
ORL 0.7234 0.7940 0.7693 0.7143 0.7529 0.7697 0.8052 0.8018 0.7813 0.8053
UMIST 0.6410 0.6190 0.6377 0.5963 0.6404 0.6469 0.6420 0.6719 0.6503 0.6791
USPS 0.6147 0.6615 0.6078 0.5865 0.6059 0.5989 0.6306 0.6768 0.5833 0.6440
Yaleface 0.4389 0.4975 0.5018 0.5099 0.4779 0.5064 0.5356 0.5382 0.5461 0.5330
@@
Table 4. Purity results on 7 benchmark datasets from 7 comparative algorithms and FKMWC with distance by Eq. (15) (Ours_bw) ,squared Euclid distance (Ours_d) and K nearest neighbors distance (Ours_kd), with the highest value represented in bold.
@@
Datasets K-Means++ Ksum Ksumx RKM CDKM MSFCM ULGE Ours_bw Ours_d Ours_kd
AR 0.2749 0.3686 0.3236 0.3215 0.2862 0.2942 0.3676 0.4071 0.3006 0.3917
JAFFE 0.7455 0.8789 0.8977 0.8310 0.7441 0.8404 0.9249 0.9155 0.8404 0.9671
MSRC_V2 0.6276 0.6910 0.6910 0.6333 0.6795 0.7381 0.7429 0.7714 0.7667 0.8095
ORL 0.5705 0.6562 0.6060 0.5200 0.6090 0.6250 0.6700 0.6900 0.6425 0.7100
UMIST 0.5110 0.4553 0.4715 0.4400 0.5043 0.5113 0.5322 0.5443 0.5270 0.5809
USPS 0.6803 0.7677 0.7240 0.6827 0.6746 0.6717 0.7530 0.7923 0.7013 0.7553
Yaleface 0.4030 0.4733 0.4915 0.4848 0.4188 0.4606 0.5212 0.5030 0.4788 0.5333
4.3. Metrics
  1. (1)
    Accuracy (ACC): It represents the clustering accuracy, defined as:
    (23) $ACC=\frac{\sum_{i=1}^{n}\delta(a_{i},\text{map}(b_{i}))}{n}$
    (24) $\delta(x,y)=\begin{cases}1,&\text{if }x=y;\\ 0,&\text{if }x\neq y.\end{cases}$
  2. (2)
    (25) $NMI(P,Q)=\frac{MI(P,Q)}{\sqrt{H(P)H(Q)}}$
  3. (3)
    (26) $\text{Purity}=\frac{n_{1}}{n_{1}+n_{2}}$
    where $n_{1}$ is the number of sample pairs correctly clustered, and $n_{2}$ is the number of incorrectly clustered pairs. All metrics range from 0 to 1, with higher values indicating better performance.
4.4. Parameter Setting
  • When using Euclidean distance:
    • $\lambda=1$ to $10$ for the AR, JAFFE, ORL, UMIST, and Yaleface datasets.
    • $\lambda=10$ to $20$ for the MSRC_V2 and USPS datasets.
  • When using the K-nearest neighbor distance:
    • $\lambda=1$ to $10$ for the JAFFE and USPS datasets.
    • $\lambda=10$ to $20$ for the MSRC_V2 dataset.
    • $\lambda=20$ to $30$ for the ORL dataset.
    • $\lambda=30$ to $40$ for the UMIST dataset.
    • $\lambda=50$ to $70$ for the Yaleface dataset.
4.5. Evaluation on Clustering Performance
Table 23, and 4 present the performance metrics of FKMWC and seven competing methods, including traditional $K$-Means, Fuzzy $K$-Means, and graph-based clustering algorithms. In cases of varying experimental results, each experiment was repeated 10 times, and the average was taken as the final result. Our variant, Ours_d, which employs squared Euclidean distance in FKMWC, consistently outperforms $K$-Means++ (also based on Euclidean distance) across all seven benchmark datasets, demonstrating both superior and more consistent performance. In comparison to RKM, Ours_d shows higher performance on the JAFFE, MSRC_V2, ORL, UMIST, and USPS datasets. Similarly, Ours_d surpasses MSFCM in most datasets. Additionally, Ours_kd, utilizing the K-nearest neighbor distance, outperforms all other comparative algorithms, including Ksum_x, which also employs K-nearest neighbor distance, as well as other graph-based algorithms like Ksum and ULGE.
4.6. Parameter Sensitivity Analysis
@@
Figure 1. Performance of the FKMWC vs. $\lambda$ on 6 benchmark datasets with squared Euclidean distance (Ours_d) and K-nearest neighbor distance (Ours_kd).
@@
Figure 2. The value of objective function (8) and clustering performance with iterations on 6 benchmark datasets.
4.7. Convergence Analysis
@@ 5. Conclusion
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.

### v2

@@ 3.2 Distance matrix
3.2.1 Squared Euclidean distance
The squared Euclidean distance measures the absolute distance between points in a Euclidean space. The distance between the $i$-th and $j$-th samples can be expressed as follows:
$d_{euc(ij)}=\left\|\textbf{x}_{i}-\textbf{x}_{j}\right\|_{2}^{2}$ (14)
3.2.2 K-Nearest Neighbor Distance
The K-nearest neighbor distance refers to the process of computing the distances between a given query point and all other points in the dataset, and then selecting the K nearest points.
$d_{knn(ij)}=\begin{cases}\|\textbf{x}_{i}-\textbf{x}_{j}\|_{2}^{2},&\textbf{x}_{i}\leftrightarrow\textbf{x}_{j}\\ \sigma,&\text{otherwise}\end{cases}$ (15)
@@ 3.2.2 K-Nearest Neighbor Distance
3.2.3 Butterworth Distance
Given an adjacency matrix $\mathbf{S}$, the distance matrix $\mathbf{D}$ can be computed utilizing a Butterworth filter as proposed in [25]. The process is outlined as follows:
$d_{bw(ij)}=\sqrt{\frac{1}{1+(\frac{s_{ij}}{\Omega})^{4}}}$ (16)
@@ 3.3 Optimization Algorithm
$J(\textbf{Y})=\textrm{tr}(\textbf{Y}^{\textrm{T}}\textbf{D}\textbf{Y}\textbf{P}^{-1})+\lambda\left\|\textbf{Y}\right\|_{F}^{2}$ (18)
@@
$\frac{\partial J}{\partial\textbf{Y}}=(\textbf{D}+\textbf{D}^{\textrm{T}})\textbf{YP}^{-1}+2\lambda\textbf{Y}+\frac{\partial(\textbf{a}^{\textrm{T}}\hat{\textbf{p}})}{\partial\textbf{Y}}$ (19)
@@
$\frac{\partial(\textbf{a}^{\textrm{T}}\hat{\textbf{p}})}{\partial y_{ij}}=\frac{\partial}{\partial y_{ij}}\left(\sum\limits_{j=1}^{K}{a_{j}p_{jj}^{-1}}\right)=-a_{j}p_{jj}^{-2}$ (20)
@@
$\frac{\partial J}{\partial y_{ij}}=g_{ij}-a_{j}p_{jj}^{-2}$ (21)
@@
$y_{ij}\leftarrow y_{ij}\sqrt{\frac{a_{j}p_{jj}^{-2}}{g_{ij}}}$ (23)
@@
7:     Normalize Y by $y_{ij}=\frac{y_{ij}}{\sum\limits_{j}{y_{ij}}}$
@@
Algorithm 1 Solving the Model (13)
3.4 Computational Complexity Analysis
@@ 4 Experiments
4.1 Benchmark Datasets
The performance of our method is evaluated on seven benchmark datasets, and the information of size is shown in Table I.
  1. 1.
    AR [26] includes 120 kinds of face with 3120 images.
  2. 2.
    JAFFE [27] consists of 213 images of different facial expressions from 10 different Japanese female subjects.
  3. 3.
    MSRC_V2 [28] which includes 7 kinds of objects with 210 images. We selected the 576-D HOG feature as the single view dataset.
  4. 4.
    ORL [29] includes 400 pictures of face from 40 people.
  5. 5.
    UMIST [30] consists of 564 images of face from 20 individuals
  6. 6.
    USPS [31] consists of 3000 handwritten digital images for each of the 10 digits from 0 to 9.
  7. 7.
    Yaleface [32] includes 165 grayscale images in GIF format of 15 people.
TABLE I: The information of the benchmark datasets, including
@@ 4.1 Benchmark Datasets
TABLE II: ACC results on 7 benchmark datasets from 7 comparative algorithms and FKMWC with different distance, with the highest value represented in bold.
@@
Datasets K-Means++ Ksum Ksumx RKM CDKM MSFCM ULGE Ours_bw Ours_d Ours_ker Ours_kd
AR 0.2514 0.2970 0.2454 0.2641 0.2653 0.2747 0.3442 0.3917 0.2843 0.2891 0.3728
JAFFE 0.7085 0.8789 0.8930 0.8310 0.7108 0.8357 0.9249 0.9155 0.8404 0.9108 0.9671
MSRC_V2 0.6052 0.7524 0.6852 0.6286 0.6657 0.7381 0.7429 0.7714 0.7667 0.7619 0.8095
ORL 0.5198 0.6337 0.5877 0.5000 0.5507 0.5725 0.6200 0.6625 0.6075 0.6250 0.6875
UMIST 0.4339 0.4209 0.4296 0.4209 0.4210 0.4591 0.5043 0.4974 0.4713 0.5113 0.5252
USPS 0.6491 0.7664 0.7240 0.6673 0.6416 0.6577 0.7530 0.7923 0.7013 0.7110 0.7553
Yaleface 0.3812 0.4339 0.4418 0.4485 0.3964 0.4545 0.5212 0.4909 0.4788 0.4909 0.5273
@@
TABLE III: NMI results on 7 benchmark datasets from 7 comparative algorithms and FKMWC with different distance, with the highest value represented in bold.
@@
Datasets K-Means++ Ksum Ksumx RKM CDKM MSFCM ULGE Ours_bw Ours_d Ours_ker Ours_kd
AR 0.5574 0.5963 0.5676 0.5752 0.5700 0.5811 0.6090 0.6500 0.5722 0.5766 0.6268
JAFFE 0.8010 0.8764 0.9013 0.8159 0.7981 0.8806 0.9005 0.9225 0.8545 0.9175 0.9623
MSRC_V2 0.5280 0.6111 0.5753 0.5612 0.5693 0.6010 0.5870 0.6265 0.6475 0.6044 0.6712
ORL 0.7234 0.7940 0.7693 0.7143 0.7529 0.7697 0.8052 0.8018 0.7813 0.7850 0.8053
UMIST 0.6410 0.6190 0.6377 0.5963 0.6404 0.6469 0.6420 0.6719 0.6503 0.6759 0.6791
USPS 0.6147 0.6615 0.6078 0.5865 0.6059 0.5989 0.6306 0.6768 0.5833 0.6192 0.6440
Yaleface 0.4389 0.4975 0.5018 0.5099 0.4779 0.5064 0.5356 0.5382 0.5461 0.5439 0.5330
@@
TABLE IV: Purity results on 7 benchmark datasets from 7 comparative algorithms and FKMWC with different distance, with the highest value represented in bold.
@@
Datasets K-Means++ Ksum Ksumx RKM CDKM MSFCM ULGE Ours_bw Ours_d Ours_ker Ours_kd
AR 0.2749 0.3686 0.3236 0.3215 0.2862 0.2942 0.3676 0.4071 0.3006 0.3080 0.3917
JAFFE 0.7455 0.8789 0.8977 0.8310 0.7441 0.8404 0.9249 0.9155 0.8404 0.9108 0.9671
MSRC_V2 0.6276 0.7524 0.6910 0.6333 0.6795 0.7381 0.7429 0.7714 0.7667 0.7619 0.8095
ORL 0.5705 0.6562 0.6060 0.5200 0.6090 0.6250 0.6700 0.6900 0.6425 0.6675 0.7100
UMIST 0.5110 0.4553 0.4715 0.4400 0.5043 0.5113 0.5322 0.5443 0.5270 0.5635 0.5809
USPS 0.6803 0.7677 0.7240 0.6827 0.6746 0.6717 0.7530 0.7923 0.7013 0.7110 0.7553
Yaleface 0.4030 0.4733 0.4915 0.4848 0.4188 0.4606 0.5212 0.5030 0.4788 0.4909 0.5333
@@
Figure 1: Performance of the FKMWC vs. $\lambda$ on 6 benchmark datasets with squared Euclidean distance (Ours_d) and K-nearest neighbor distance (Ours_kd).
@@
Figure 2: The value of objective function (13) and clustering performance with iterations on 6 benchmark datasets.
4.2 Comparison Algorithms
  1. 1.
    K-Means++ [14]:The advantages of careful seeding for K-Means.
  2. 2.
    Ksum [33]: This algorithm integrates spectral clustering and K-Means into a unified framework.
  3. 3.
    Ksum-x [33]: Distinct from Ksum, Ksum-x inputs a data matrix instead of a K-NN graph.
  4. 4.
    RKM [34]: This algorithm minimizes the squared distance from the samples to the clustering centers and incorporates a regularization term for balanced clustering.
  5. 5.
    CDKM [35]: Demonstrates that the underlying model of K-Means can be transformed into a matrix trace maximization problem, proposing an efficient coordinate descent solution.
  6. 6.
    MSFCM [36]A new membership scaling Fuzzy C-Means clustering algorithm.
  7. 7.
    ULGE [37] Unsupervised Large Graph Embedding.
4.3 Metrics
  1. 1.
    Accuracy (ACC): It represents the clustering accuracy, defined as:
    $ACC=\frac{\sum_{i=1}^{n}\delta(a_{i},\text{map}(b_{i}))}{n}$ (24)
    $\delta(x,y)=\begin{cases}1,&\text{if }x=y;\\ 0,&\text{if }x\neq y.\end{cases}$ (25)
  2. 2.
    $NMI(P,Q)=\frac{MI(P,Q)}{\sqrt{H(P)H(Q)}}$ (26)
  3. 3.
    $\text{Purity}=\frac{n_{1}}{n_{1}+n_{2}}$ (27)
    where $n_{1}$ is the number of sample pairs correctly clustered, and $n_{2}$ is the number of incorrectly clustered pairs.
    All metrics range from 0 to 1, with higher values indicating better performance.
4.4 Parameter Setting
  • When using Euclidean distance:
    • $\lambda=1$ to $10$ for the AR, JAFFE, ORL, UMIST, and Yaleface datasets.
    • $\lambda=10$ to $20$ for the MSRC_V2 and USPS datasets.
  • When using the K-nearest neighbor distance:
    • $\lambda=1$ to $10$ for the JAFFE and USPS datasets.
    • $\lambda=10$ to $20$ for the MSRC_V2 dataset.
    • $\lambda=20$ to $30$ for the ORL dataset.
    • $\lambda=30$ to $40$ for the UMIST dataset.
    • $\lambda=50$ to $70$ for the Yaleface dataset.
4.5 Evaluation on Clustering Performance
Table IIIII, and IV present the performance metrics of FKMWC and seven competing methods, including traditional $K$-Means, Fuzzy $K$-Means, and graph-based clustering algorithms. In cases of varying experimental results, each experiment was repeated 10 times, and the average was taken as the final result. Our variant, Ours_d, which employs squared Euclidean distance in FKMWC, consistently outperforms $K$-Means++ (also based on Euclidean distance) across all seven benchmark datasets, demonstrating both superior and more consistent performance. In comparison to RKM, Ours_d shows higher performance on the JAFFE, MSRC_V2, ORL, UMIST, and USPS datasets. Similarly, Ours_d surpasses MSFCM in most datasets. Additionally, Ours_kd, utilizing the K-nearest neighbor distance, outperforms all other comparative algorithms, including Ksum_x, which also employs K-nearest neighbor distance, as well as other graph-based algorithms like Ksum and ULGE. Ours_ker, which employs kernel distance in FKMWC, due to its calculation of the distance in the kernel space, demonstrates good applicability for linearly inseparable datasets, and achieves higher results than squared Euclidean distance on six datasets. It only achieves similar results as Euclidean distance on the MSRC_V2 dataset.
4.6 Parameter Sensitivity Analysis
4.7 Convergence Analysis